OpenStudy (anonymous):

1 and 1/3 times 5 and 5/8.

OpenStudy (anonymous):

this is the same as (4/3) * (45/8) = (4*45)/(3*8) = (180/24) = 14/2

OpenStudy (anonymous):

\[1\tfrac{1}{3} \times 5\tfrac{5}{8}=\frac{4}{3}\times \frac{45}{8}\] and go from there. some cancellation to do first

OpenStudy (anonymous):

i think you can do this - plug it into your calculator like this (1+(1/3))(5+(5/8))

OpenStudy (mathteacher1729):

Numerator = top Denominator = bottom To ADD fractions, you need them to have the same denominator. To MULTIPLY fractions, just multiply the numerators times the numerators and the denominators time the denominators. If you have the same number in the numerator and the denominator, that is equal to one. So "1 and 1/3" is "3/3 + 1/3" Which (by the Adding rule) is 4/3. "5 and 5/8" is "5 times 8 all over 8 + 5/8" or "40/8 +5/8" which is "45/8". Now multiply 4/3 times 45/8. :) That's how you get the answer.

OpenStudy (anonymous):

Also if you find commonnality (lol made up word i believe) between the denominators and the numerators such as your problem has a 4 and and an 8 you can get rid of the 4 and turn the 8 into a 2

OpenStudy (anonymous):

and the 3 and 45 can turn into 1 and 15 so you should get 15/2